# Active Learning for Continual Learning: Keeping the Past Alive in the Present

### v1

@@ 5.3 Ablation Studies
Effect of $\lambda$. Figure 3 illustrates the influence of $\lambda$ on the fourth task of SplitCIFAR10, trained with ER, which is the parameter that dictates the quality of data to prioritize during query selection. The different values of $\lambda$ are selected by our theoretical value in Theorem 4.1 for different tasks in SplitCIFAR10. For a fair comparison, we choose identical initial points for the first round of training, select subsets for various $\lambda$ values, and then re-train it to examine the impact of a change in $\lambda$. It is observed that the forgetting rises when $\lambda$ is small because AccuACL focuses on the new task, and vice versa. Furthermore, we can see that the accuracy reaches the optimal at the theoretical value of $\lambda$.
@@
Scoring Methods. Table 5.3 shows the effect of different scoring methods defined in Section 4.5. Querying based solely on $\mathcal{M}(\cdot)$ focuses exclusively on learning new information, neglecting the target Fisher information defined in Section 4.2. This approach leads to high forgetting and low accuracy. In contrast, querying based solely on $\mathcal{D}(\cdot)$ emphasizes the distribution of information across parameters. While this approach selects examples that balance information between past and new tasks, it neglects the amount of each example’s informativeness. As a result, although it may reduce forgetting, it does not greatly enhance average accuracy. By considering both scores together, AccuACL achieves the highest average accuracy with maintaining low forgetting. Moreover, employing $\mathcal{D}(\cdot)$ as the primary criterion for selection allows for an AL algorithm that is more aligned with the motivation of our paper, which is to maintain a balance between the two learning properties of ACL: preventing catastrophic forgetting, and quick learning of new tasks. Ensuring the magnitude of informativeness after establishing the balance has shown effectiveness through empirical evaluation.
Previous AL with Memory. To show how effectively AccuACL uses the information of past data via a memory buffer, we run experiments on the greedy algorithms, kCenter and BADGE, to use the examples in the memory buffer as an initial selected subset to incorporate the past knowledge into the algorithms. Surprisingly, as in Table 2, using the memory buffer as a starting point even degrades the performance of the original AL algorithms. We conjecture that additionally using the memory buffer may force AL algorithms to concentrate on learning new information even more as they attempt to choose the examples that are the most dissimilar from past data.
Evaluating Learning Accuracy in ACL. Learning accuracy Mirzadeh et al. (2022) is an essential evaluation for assessing the plasticity of CL systems. Within the context of ACL, it indicates the efficacy of AL algorithms in quickly adapting to new tasks. To investigate this, we performed an additional experiment evaluating the learning accuracy of AL strategies on the SplitCIFAR100 dataset($M$=500), optimized by DER++. As can be seen in Table 3, traditional AL baselines achieve superior learning accuracy, consistent with our motivation that traditional AL approaches emphasize fast adaptation to new tasks. On the other hand, AccuACL surpasses the AL baselines in forgetting, as AccuACL specifically tackles another essential perspective of ACL learning properties: the prevention of catastrophic forgetting. By successfully balancing these two learning properties, AccuACL exhibits state-of-the-art performance in average accuracy, highlighting its value in overall ACL performance. Moreover, as AccuACL selects examples that mitigate recency bias in linear classifiers Mai et al. (2021), it allows reliable Fisher information matrix calculation for the upcoming AL rounds.
@@ 6 Conclusion
Appendix E Extended Experiments
E.1 Evaluating the Impact of AL on Inter-Task Discrimination-based CL
Table 5: Performance comparison of AL baselines and AccuACL combined with LODE (Liang and Li, 2024) on SplitCIFAR10, SplitCIFAR100, and SplitTinyImageNet.
AL Method SplitCIFAR10 SplitCIFAR100 SplitTinyImageNet
M=100 M=500 M=2000
$A_{5}(\uparrow)$ $F_{5}(\downarrow)$ $A_{10}(\uparrow)$ $F_{10}(\downarrow)$ $A_{10}(\uparrow)$ $F_{10}(\downarrow)$
Uniform 37.6
$\pm$0.8
25.2
$\pm$5.9
27.4
$\pm$1.0
28.4
@@ E.1 Evaluating the Impact of AL on Inter-Task Discrimination-based CL
29.0
$\pm$0.3
11.2
$\pm$0.2
24.2
$\pm$0.6
AccuACL 36.4
$\pm$1.0
24.2 $\pm$1.0 28.4 $\pm$0.5 26.4 $\pm$0.7 12.4 $\pm$0.9 23.5
$\pm$0.6
To demonstrate that AL strategies addressing catastrophic forgetting are advantageous even in CL methods focusing on inter-task discrimination (Wang et al., 2022a, Lin et al., 2023), we conducted additional experiments on LODE (Liang and Li, 2024). LODE decouples the loss into intra-task and inter-task distinction, using the hyperparameter $\rho$ to control the degree of inter-task distinction. As shown in Table 5, AccuACL consistently outperforms the AL baselines across the majority of settings together. Entropy and LeastConf exhibited low forgetting on SplitTinyImageNet, attributable to the exceedingly low average accuracy. Moreover as shown in Figure 4, AccuACL exhibits superior performance across various selections of $\rho$, i.e., varied degrees of inter-task discrimination.
Figure 4: Performance comparison of AccuACL and baseline methods on SplitCIFAR100($M$=500) using LODE (Liang and Li, 2024) across different values of $\rho$.
E.2 Performance Analysis of AL strategies Throughout AL Rounds
@@
Figures 5(a) and 5(b) show the performance of ER over the AL rounds at the 5-th task on SplitCIFAR100. Here, AccuACL performs the best, and the performance gap to the AL baselines gets larger as the AL rounds progress. Notably, AccuACL shows very low forgetting, indicating its superiority in preventing catastrophic forgetting. Interestingly, the performances peak around the tenth round of AL. Since the AL was conducted up to 10 rounds for previous tasks, extending beyond 10 rounds for the new task results in an imbalanced data distribution, likely causing the observed performance drop.
E.3 The Efficacy of Fisher Information Matrix Estimation via Memory Buffer
@@
To investigate the influence of using a memory buffer for estimating the target Fisher information matrix in rehearsal-based CL, we designed a modified version of AccuACL, AccuACLFull. In contrast to the original AccuACL, which employs a memory buffer, AccuACLFull leverages all data from previously seen tasks to approximate the target Fisher information matrix. As shown in Table 6, AccuACL achieves competitive results compared to AccuACLFull, indicating that a minimal memory buffer(1K examples) suffices for adequate estimation. While AccuACLFull reduces forgetting due to more comprehensive parameter estimation, it results in lower average accuracy. We hypothesize that AccuACLFull’s extensive data use of the past improves its capacity to retain important parameters from previous tasks that cannot be effectively captured by replaying a memory buffer. However, in a practical scenario of rehearsal-based CL, estimating the Fisher information matrix using only the memory buffer strikes a better balance between retaining past knowledge and integrating new tasks by efficiently selecting new task samples without conflicts with the memory data, instead of attempting to preserve every knowledge of the past.
E.4 Robustness of AccuACL with Varying Task Orders
Table 7: Performance comparison of AccuACL and Uniform across multiple task order permutation on SplitCIFAR100($M$=500), trained with ER.
AL Method Sequential Perm. #1 Perm. #2 Perm. #3 Perm. #4
Uniform 10.9
$\pm$0.4
63.7
$\pm$0.6
11.7
$\pm$0.4
62.0
$\pm$0.3
11.1
$\pm$0.2
61.0
$\pm$1.2
11.9
$\pm$0.4
62.5
$\pm$0.5
11.9
$\pm$0.4
61.4
$\pm$0.5
To further verify the robustness of our technique, we ran additional experiments on the SplitCIFAR100($M$=500) dataset, where the original data is split into ten tasks of ten classes each in sequential order. To investigate the impact of task order, we generated four random permutations of the tasks. Table 7 demonstrates that AccuACL consistently outperforms the second-best method, Uniform, in all permutations. Our results indicate that our method is highly robust to changes in task order, further establishing its efficacy.
Appendix F Visualization of AL Query
Figure 6: Query examples for each task in SplitCIFAR10, for four AL methods: Entropy, kCenter, BADGE, and AccuACL.
To provide more insights into AccuACL regarding its visual features, we demonstrate query examples for each task in SplitCIFAR10, corresponding to four AL methods in Figure 6. The sampling process is conducted from each checkpoint model of ER trained on SplitCIFAR10($M$=100).
Appendix G Implementation Details
Table 8: Detailed experiment settings.
Hyperparamters SplitCIFAR10 SplitCIFAR100 SplitTinyImageNet
Training Configuration architecture ResNet18 ResNet18 ResNet18
training epoch 50 20 30
batch size 16 16 32
optimizer SGD SGD SGD
momentum 0.8 0.8 0.9
weight decay $10^{-4}$ $10^{-4}$ 0
learning rate (lr) 0.01 0.01 0.01
lr scheduler Cosine Annealing Cosine Annealing Cosine Annealing
For each CL task, we perform multi-round ACL following the conventional AL setup (Ash et al., 2019). For each experiment, we use two different sizes for a memory buffer, $\{100,200\}$ for SplitCIFAR10, $\{500,1000\}$ for SplitCIFAR100, and $\{2000,5000\}$ for SplitTinyImageNet. We query 1000, 2000, and 3000 examples during ten rounds of AL for SplitCIFAR10, SplitCIFAR100, and SplitTinyImageNet for each task, respectively. In terms of over-sampling in Section 4.5, we oversample two times the original query size.
The overall optimization and training setup are shown in Table 8. Among the four CL algorithms, GSS (Aljundi et al., 2019) and DER (Buzzega et al., 2020) require hyperparameters to fix. For GSS, the number of random samples for determining the maximal cosine similarity is set to 5. For DER, the values of $\alpha$ and $\beta$ are given weights of 0.1 and 0.5, respectively. $\alpha$ represents the weight allocated to the mean-squared error, while $\beta$ represents the weight assigned to the cross-entropy error. All AL baselines, except for BAIT (Ash et al., 2021), need no additional hyperparameters to be specified. The oversampling rate for forward greedy optimization is configured as 2. For AccuACL, we choose the top-10 embeddings per class for SplitCIFAR10 and the top-5 embeddings for SplitCIFAR100 and SplitTinyImageNet in order to compute the distribution score. We conduct our experiment based on the Avalanche codebase (Carta et al., 2023). All of our experiments are implemented with PyTorch 1.12.1 and performed with NVIDIA GeForce RTX 4090 24GB on CUDA version 12.0. All experiments except for BAIT are repeated three times, and the average and standard deviation are reported.
Appendix H Limitations & Future Works
@@
While AccuACL has repeatedly shown its efficacy in ACL scenarios, we have yet to validate the suitability of our approach in realistic CL situations, such as noisy labels Bang et al. (2022), Park et al. (2022), blurry tasks Bang et al. (2021), and imbalanced data Chrysakis and Moens (2020). Moreover, as AccuACL necessitates the use of memory buffer from rehearsal-based CL, a rehearsal-free approach of AccuACL will be a promising approach, possibly with the use of regularization-based CL methods to further accurately maintain the target Fisher information matrix without a memory buffer. Furthermore, as AL for vision-language model (VLM) Bang et al. (2024) has shown possibilities, devising an ACL strategy for VLM to can be a promising research field.

### v2

@@ 5.3 More Results and Ablation Studies
Effect of $\lambda$. Figure 3 illustrates the influence of $\lambda$ on the fourth task of SplitCIFAR10, trained with ER, which is the parameter that dictates the quality of data to prioritize during query selection. The different values of $\lambda$ are selected by our theoretical value in Theorem 4.1 for different tasks in SplitCIFAR10. For a fair comparison, we choose identical initial points for the first round of training, select subsets for various $\lambda$ values, and then re-train it to examine the impact of a change in $\lambda$. It is observed that the forgetting rises when $\lambda$ is small because AccuACL focuses on the new task, and vice versa. Furthermore, the accuracy reaches the optimal at the theoretical value of $\lambda$.
AL Baselines with Memory. To show how effectively AccuACL uses the information of past data via a memory buffer, we run experiments on the greedy algorithms, kCenter and BADGE, to use the examples in the memory buffer as an initially selected subset to incorporate the past knowledge into the algorithms. Surprisingly, as in Table 2, using the memory buffer as a starting point even degrades the performance of the original AL algorithms. We conjecture that additionally using the memory buffer may force AL algorithms to concentrate on learning new information even more as they attempt to choose the examples that are the most dissimilar from past data. This finding indicates that simply providing AL with past data is not sufficient, whereas AccuACL properly balances past and new knowledge through data-driven information. For further analysis on the impact of using a memory buffer in AccuACL, please refer to Appendix F.3.
Learning Accuracy in ACL. Learning accuracy (Mirzadeh et al., 2022) is an essential evaluation for assessing the plasticity of CL algorithms, which is the accuracy of the new task immediately after learning. Within the context of ACL, it indicates the efficacy of AL algorithms in quickly adapting to new tasks. Accrodingly, we perform an additional experiment evaluating the learning accuracy of AL strategies on the SplitCIFAR100 dataset($M$=500), optimized by DER++. As can be seen in Table 3, traditional AL baselines achieve superior learning accuracy, consistent with our motivation that traditional AL approaches emphasize fast adaptation to new tasks. On the other hand, AccuACL surpasses the AL baselines in forgetting, as AccuACL specifically tackles another essential perspective of ACL learning properties: prevention of catastrophic forgetting. By successfully balancing these two learning properties, AccuACL exhibits state-of-the-art performance in average accuracy, highlighting its value in overall ACL performance. Moreover, as AccuACL selects examples that mitigate recency bias in linear classifiers (Mai et al., 2021), it allows reliable Fisher information matrix calculation for the upcoming AL rounds.
@@
Scoring Methods. Table 5.3 shows the effect of different scoring methods defined in Section 4.5. Querying based solely on $\mathcal{M}(\cdot)$ focuses exclusively on learning new information, neglecting the target Fisher information defined in Section 4.2. This approach leads to high forgetting and low accuracy. In contrast, querying based solely on $\mathcal{D}(\cdot)$ emphasizes the distribution of information across parameters. While this approach selects examples that balance information between past and new tasks, it neglects the amount of each example’s informativeness. As a result, although it may reduce forgetting, it does not greatly enhance average accuracy. AccuACL employs $\mathcal{D}(\cdot)$ as the primary selection criterion for querying, aligning with the objective of our study to sustain a balance between the two learning properties of ACL—preventing catastrophic forgetting and quickly learning new tasks. Ensuring the magnitude of informativeness after establishing the balance has shown effectiveness through empirical evaluation.
@@ 6 Conclusion
Acknowledgments
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II200862, DB4DL: High-Usability and Performance In-Memory Distributed DBMS for Deep Learning, 50% and No. RS-2022-II220157, Robust, Fair, Extensible Data-Centric Continual Learning, 50%). Jae-Gil Lee was partly supported by Samsung Electronics Co., Ltd. (IO201211-08051-01).
@@ Acknowledgments
Appendix E Implementation Details
Table 5: Detailed experiment settings.
Hyperparamters SplitCIFAR10 SplitCIFAR100 SplitTinyImageNet
Training Configuration architecture ResNet18 ResNet18 ResNet18
training epoch 50 20 30
batch size 16 16 32
optimizer SGD SGD SGD
momentum 0.8 0.8 0.9
weight decay $10^{-4}$ $10^{-4}$ 0
learning rate (lr) 0.01 0.01 0.01
lr scheduler Cosine Annealing Cosine Annealing Cosine Annealing
For each CL task, we perform multi-round ACL following the conventional AL setup (Ash et al., 2019). For each experiment, we use two different sizes for a memory buffer, $\{100,200\}$ for SplitCIFAR10, $\{500,1000\}$ for SplitCIFAR100, and $\{2000,5000\}$ for SplitTinyImageNet. We query 1000, 2000, and 3000 examples during ten rounds of AL for SplitCIFAR10, SplitCIFAR100, and SplitTinyImageNet for each task, respectively. In terms of over-sampling in Section 4.5, we over-sample two times the original query size.
The overall optimization and training setups are shown in Table 5. Among the four CL algorithms, GSS (Aljundi et al., 2019) and DER (Buzzega et al., 2020) require hyperparameters to fix. For GSS, the number of random samples for determining the maximal cosine similarity is set to 5. For DER, the values of $\alpha$ and $\beta$ are given weights of 0.1 and 0.5, respectively. $\alpha$ represents the weight allocated to the mean-squared error, while $\beta$ represents the weight assigned to the cross-entropy error. All AL baselines, except for BAIT (Ash et al., 2021), need no additional hyperparameters to be specified. The oversampling rate for forward greedy optimization is configured as 2. For AccuACL, we choose the top-10 embeddings per class for SplitCIFAR10 and the top-5 embeddings for SplitCIFAR100 and SplitTinyImageNet in order to compute the distribution score. We conduct our experiment based on the Avalanche codebase (Carta et al., 2023). All experiments are implemented with PyTorch 1.12.1 and performed with NVIDIA GeForce RTX 4090 24GB on CUDA version 12.0. All experiments except for BAIT are repeated three times, and the average and standard deviation are reported.
Appendix F Extended Experiments
F.1 Robustness of AccuACL with Varying Task Orders
Table 6: Performance comparison of AccuACL and Uniform across multiple task order permutations on SplitCIFAR100($M$=500), trained with ER.
AL Method Sequential Perm. #1 Perm. #2 Perm. #3 Perm. #4
Uniform 10.9
$\pm$0.4
63.7
$\pm$0.6
11.7
$\pm$0.4
62.0
$\pm$0.3
11.1
$\pm$0.2
61.0
@@ F.1 Robustness of AccuACL with Varying Task Orders
To further verify the robustness of our technique, we run additional experiments on the SplitCIFAR100($M$=500) dataset, where the original data is split into ten tasks of ten classes each in sequential order. To investigate the impact of task order, we generate four random permutations of the tasks. Table 6 demonstrates that AccuACL consistently outperforms the second-best method, Uniform, in all permutations. Our results indicate that our method is highly robust to changes in task order, further establishing its efficacy.
F.2 Performance Analysis of AL Strategies throughout AL Rounds
@@
Figures 4(a) and 4(b) show the performance of ER over the AL rounds at the fifth task on SplitCIFAR100. Here, AccuACL performs the best, and the performance gap to the AL baselines gets larger as the AL rounds progress. Notably, AccuACL shows very low forgetting, indicating its superiority in preventing catastrophic forgetting. Interestingly, the performances peak around the tenth round of AL. Since the AL was conducted up to 10 rounds for previous tasks, extending beyond 10 rounds for the new task results in an imbalanced data distribution, likely causing the observed performance drop.
F.3 Efficacy of Fisher Information Matrix Estimation via Memory Buffer
@@
To investigate the influence of using a memory buffer for estimating the target Fisher information matrix in rehearsal-based CL, we design a modified version of AccuACL, AccuACLFull. In contrast to the original AccuACL, which employs a memory buffer, AccuACLFull leverages all data from previously seen tasks to approximate the target Fisher information matrix. As shown in Table 7, AccuACL achieves competitive results compared to AccuACLFull, indicating that a minimal memory buffer (1000 examples) suffices for adequate estimation. While AccuACLFull reduces forgetting due to more comprehensive parameter estimation, it results in lower average accuracy. We hypothesize that AccuACLFull’s extensive data use of the past improves its capacity to retain important parameters from previous tasks that cannot be effectively captured by replaying a memory buffer. However, in a practical scenario of rehearsal-based CL, estimating the Fisher information matrix using only the memory buffer strikes a better balance between retaining past knowledge and integrating new tasks by efficiently selecting new task examples without conflicts with the memory data, instead of attempting to preserve every knowledge of the past.
Appendix G Limitations & Future Works
@@
While AccuACL has repeatedly shown its efficacy in ACL scenarios, we have yet to validate the suitability of our approach in realistic CL situations, such as noisy labels (Bang et al., 2022, Park et al., 2022), blurry tasks (Bang et al., 2021), and imbalanced data (Chrysakis and Moens, 2020). Moreover, as AccuACL necessitates the use of memory buffer from rehearsal-based CL, a rehearsal-free approach of AccuACL will be a promising approach, possibly with the use of regularization-based CL methods to further accurately maintain the target Fisher information matrix without a memory buffer. Furthermore, as AL for vision-language models (VLMs) (Bang et al., 2024) has shown possibilities, devising an ACL strategy for VLMs to can be a promising research field.